OpenStudy (anonymous):

((x^2)+4x/x-5)>0 What is the set of solutions to this problem?

OpenStudy (anonymous):

is that x-5 the denominator like this? \[\large x^2+\frac{4x}{x-5} \]>0

OpenStudy (anonymous):

no, (x^2)+4x is the numerator. and then x-5 is the denominator.

OpenStudy (anonymous):

so it's like this... \[\large \frac{x^2+4x}{x-5}>0 \]

OpenStudy (anonymous):

yup ^^

OpenStudy (anonymous):

first we need to solve this... \[\large \frac{x^2+4x}{x-5}=0 \]

OpenStudy (anonymous):

so we need to factor the top... \[\large \frac{x(x+4)}{x-5}=0 \] this implies x(x-4)=0 so x= 0, x=-4 correct?

OpenStudy (anonymous):

sorry i mean x(x+4)=0 so x=0, x=-4

OpenStudy (anonymous):

these are the zeros (where the graph crosses the x axis) of the expression on the left hand side.

OpenStudy (anonymous):

yeah, i got that too. but then the question asks us to state the solution using interval notation. so will it be like: (-infinity,-4)U(0,infinity)? cuz i tried that.. and it was wrong ><

OpenStudy (anonymous):

yes... but we need to consider 1 more thing... that is the vertical asymptote... to get the vertical asymptote, you need to set the denominator equal to zero... then solve for x. what x value will make the denominator equal to zero?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

ok good... now we must check intervals...

OpenStudy (anonymous):